In maize (corn), assume that the genes A and B are linked and 30 map units apart. If a plant of Ab/aB is selfed, what proportion of the progeny would be expected to be of ab/ab genotype?

a.

30%

b.

4.5%

c.

9%

d.

15%

e.

2.25%
Answer #1

Ans - (e) 2.25%

The distance between two genes is equal to the frequency of the recombination. The distance between A and B is 30 map unit which means 30% of the progeny will be recombinant.

Ab/aB - will produce four types of gamestes

Ab and aB are parental type (non recombinant) - 70%

ab and AB are recombinant type - 30%

Frequency of the ab gamete = 30/2 = 15%

Proportion of the progeny will have ab/ab genotype =

(15/100)(15/100) *100 = 2.25%